# BEVPlace++: Fast, Robust, and Lightweight LiDAR Global Localization for Unmanned Ground Vehicles

### v2

@@ III BEV-based Global Localization Pipeline
Our method uses bird’s eye view (BEV) images as an intermediate representation of LiDAR data to perform global localization. As shown in Fig. 2, we project point clouds into BEV views and extract distinct local and global descriptors with the rotation equivariant and invariant network (REIN). Then, we employ a two-stage localization paradigm, i.e., place recognition followed by pose estimation. In place recognition, we perform place retrieval with the global descriptors and find the top-matched BEV image from a pre-built database. In pose estimation, we exploit the local descriptors to estimate the relative pose between the query and matched BEV image pair using RANSAC. The global pose of the query is finally computed as the combination of the stored pose of the matched frame and the relative pose. In the following, we first explain the BEV image generation process, and then briefly describe the proposed REIN network. Finally, we elaborate on our proposed two-stage localization inference pipeline.
@@ III-A Bird’s Eye View Representation
Following existing 3-DoF localization works [35, 16, 56], we assume that when the ground vehicle moves within a local area, it travels on a rough plane. Based on this assumption, we generate BEV images through the orthogonal projection and concentrate on estimating the pose in 3-DoF, including (x, y, yaw). The (z, pitch, roll) can also be derived from the stored pose of the matched frame, but they are not the primary focus of this article. The BEV image representation shows benefits in many aspects of our hierarchical localization system. In place recognition, BEV images offer a comprehensive view of the distribution of road elements. Therefore, it is more intuitive and stable to extract global descriptors from such images to depict the structural information of the scene. In the context of pose estimation, the transformations of BEV image pairs are estimated by solving the BEV image matching problem, which is fast because of the lightweight BEV representation. The lightweight BEV representation also benefits the storage resource which is important for real-world applications and multi-robot communication.
@@
where $N_{g}$ denotes the number of points in the grid at position $(u,v)$ and $N_{m}$ the normalization factor. $N_{m}$ is set as the max value of the point cloud density.
Unlike traditional BEV image, also called elevation map [34, 35], that stores the maximum height of the points in each bin, we use the point densities. This is because the elevation map is sensitive to the orientation of the sensor, as the maximum height recorded varies with the distance between the scanner and the object. The density of points scanned on a surface, on the other hand, is less sensitive to viewpoint changes [52, 57].
@@ III-C Place Recognition
where $\mathbf{T}_{i}\in\mathbb{R}^{3\times 3}$ and $\mathbf{V}_{i}\in\mathbb{R}^{K\times C}$ are the corresponding global pose and global descriptor of the BEV image $\mathbf{I}_{i}$. The database construction can be performed online or offline according to specific tasks. For localization within a global map, the database is typically constructed offline during the map-building process, far from the time of current application. For the task of loop closure in SLAM, the database is built in real-time, and its elements grow incrementally with the explored area of the vehicle.
2) Place retrieval. For the query BEV image $\textbf{I}_{q}$, we generate its global descriptor $\textbf{V}_{q}$ with BEVPlace++ feature extractor. The matched frame is computed through the nearest global descriptor searching:
@@ III-D Pose Estimation
1) BEV image matching. We first extract FAST [58] keypoints from the BEV images to enable fast and accurate keypoint detection. Furthermore, on BEV images, the detected FAST keypoints usually have good repeatability since they are usually located at objects with vertical structures in the environment (e.g., poles, facades, guideposts). We then assign each keypoint with a local descriptor interpolated from the REM feature map. We perform local feature matching between the BEV image pair and use RANSAC to estimate the relative transform with the keypoint correspondences.
@@ IV-A BEV Features Generation Through CNNs
We further plot the average feature distance with respect to translation displacements using different CNN backbones, including ResNet34 [26], VGG16 [59] and EfficientNet [60], on different datasets, such as KITTI [61], NCLT [62], and MCD [63], shown in Fig. 3 (e), (f), and (g). These datasets include point clouds collected by various types of LiDAR scanners with different fields of view and sparsity levels across diverse environments. As can be seen, spatially close features exhibit small feature distances while distant ones show larger feature distances, regardless of the dataset setup or used CNN backbones, which could formulated as
@@
This inherent distinctiveness reveals that deep BEV features can capture and represent unique patterns in the BEV image. Such properties are crucial for accurately estimating pose, as they help in distinguishing between different keypoints and ensuring that their spatial relationships are preserved.
@@
Leveraging this characteristic, we establish BEV feature correspondences through nearest neighbor search and utilize RANSAC to solve for pose estimation. Because our matching approach does not rely on specific dataset fine-tuning, our method can effectively adapt to point clouds from various LiDAR scanners and generalize robustly across diverse environmental conditions.
Fig. 4 (a) illustrates feature matching results using ResNet34 under large translations. Each row corresponds to BEV images from the KITTI, NCLT, and MCD datasets, respectively. Following pose estimation, the overlaid BEV image pairs indicate our method’s capability to accurately estimate significant translation movements. It is noted that feature matching could fail under rotations, as shown in Fig. 4 (b), since CNNs are not rotation-equivariant and the feature distinctiveness is only preserved under translation transformations.
@@ IV-B Rotation Equivariant Module
As discussed above, CNNs are translation equivariant but cannot handle rotations well. To achieve more robust pose estimations, we further design a network to extract rotation-equivariant feature maps from BEV images. We introduce a simple and effective rotation equivariant module (REM) to extract rotation equivariant local features. REM uses modern CNNs as basic modules to keep their representation ability to BEV images and achieves rotation equivariance by using the maximum CNN feature response of BEV images under different rotation transforms.
@@
Figure 6: Comparison between ResNet features and REM features with rotations. Features are extracted from a pair of rotated BEV images with ResNet and REM, respectively. The difference between the ResNet features of the rotated BEV and the rotated ResNet features of the original BEV is large. On the other hand, REM shows rotation equivariance, and the difference between the REM features of the rotated BEV and the rotated REM features of the original BEV tends to be zero.
@@
We apply an identical transformation $\textbf{R}_{\alpha}\textbf{R}^{-1}_{\alpha}$ to the feature map and have
$\textbf{F}^{\prime}$ $=\max_{r\in\mathcal{R}}\textbf{R}_{\alpha}\textbf{R}_{\alpha}^{-1}\textbf{R}^{-1}_{r}\circ{\phi((\textbf{R}_{r}\textbf{R}_{\alpha})\circ\textbf{I})}$ (13)
$=\textbf{R}_{\alpha}\circ\max_{r\in\mathcal{R}}\textbf{R}_{\alpha}^{-1}\textbf{R}^{-1}_{r}\circ{\phi((\textbf{R}_{r}\textbf{R}_{\alpha})\circ\textbf{I})}.$
$=\textbf{R}_{\alpha}\circ\max_{r\in\mathcal{R}}\textbf{R}_{r+\alpha}^{-1}\circ{\phi((\textbf{R}_{r+\alpha})\circ\textbf{I})}$
@@
This validates the REM feature map $\textbf{F}^{\prime}$ of the rotated image is equal to the rotated feature map $\textbf{F}_{\alpha}$ of the raw BEV image. Theoretically, we need to sample infinite angles to achieve continuous rotation equivariance. In our experiments, we find a small $N$ can achieve satisfactory performance. This is possible because the pooling and downsample operations in CNN could provide some robustness to small movements.
@@
$\mathbf{V}_{k}=\sum_{i}v_{k}(\mathbf{f}_{i})=\sum_{i}\frac{e^{\mathbf{w}_{k}^{T}\mathbf{f}_{i}+b_{k}}}{\sum_{k^{\prime}}e^{{\mathbf{w}_{k^{\prime}}^{T}\mathbf{f}_{i}+b_{k^{\prime}}}}}(\mathbf{f}_{i}-\mathbf{c}_{k}),$ (15)
where $\mathbf{w}_{k}$ and $b_{k}$ are the learnable weights and bias.
2) Rotation Invariance Analysis. Let $\{\mathbf{f}_{i}\}_{i=1,...,{H^{\prime}W^{\prime}}}$ and $\{\mathbf{f^{\prime}}_{i}\}_{i=1,...,{H^{\prime}W^{\prime}}}$ be the local descriptor sets of the raw BEV image and the rotated image. As the REM feature map is rotation equivariant, set $\{\mathbf{f}_{i}\}_{i=1,...,H^{\prime}W^{\prime}}$ and $\{\mathbf{f^{\prime}}_{i}\}_{i=1,...,H^{\prime}W^{\prime}}$ has same elements but different permutation. We have
$\mathbf{V}_{k}^{\prime}$ $=v_{k}(\mathbf{f}_{1}^{\prime})+v_{k}(\mathbf{f}_{2}^{\prime})+...+v_{k}(\mathbf{f}_{n}^{\prime})$ (16)
$=\sum_{i}v_{k}(\mathbf{f}_{i}^{\prime})$
$=\sum_{i}v_{k}(\mathbf{f}_{i})=\mathbf{V}_{k},$
indicating the rotation invariance of global descriptors.
@@
Dataset KITTI [61] NCLT  [62] MCD [63] RobotCar [65] Inhouse [14]
LiDAR Velodyne HDL-64E Velodyne HDL-32E Livox Mid-70 SICK LD-MRS 3D Velodyne HDL-64
Field of View (HxV) ${}^{\circ}\times$ 26.8 ${}^{\circ}\times$ 41.3 ${}^{\circ}\times$ 70 (non repeat) ${}^{\circ}\times$ 3.2 ${}^{\circ}\times$ 26.8
No. of Sequences 5 7 6 44 15
Scenes City, Country Campus City City Campus
Cities Karlsruhe Michigan Singapore Oxford Singapore
Trajectory length 5 km 10 km 2km 10 km 10 km
Time span Single days Across 1 year Single days Across 1 year Several days
Evaluation PR & LC & GL PR & LC PR & GL PR PR
@@
Figure 7: Sample data of different datasets, including the first-row LiDAR scanners, the second-row point clouds, and the third-row corresponding BEV images.
@@ V-A Datasets
We selected datasets to thoroughly evaluate the performance of our method in large-scale environments, under long-term changes, and with various LiDAR sensor setups. The evaluation datasets include KITTI [61], NCLT [62], MCD [63], the RobotCar [65], and In-house dataset [14], which were collected in different cities.
We evaluate loop closing performance on KITTI and NCLT datasets since their data sequences have large loops. We test global localization on KITTI, NCLT, and MCD datasets as they have accurate ground truth poses. We evaluate place recognition on all the datasets. Our evaluation datasets cover diverse scenes, including city, countryside, and campus, and are collected in large-scale places under large time spans. The point cloud data are of different sparsity and different fields of view due to the usage of the various types of LiDAR sensors. Tab. I summarizes the meta information of the datasets. Fig. 7 provides some example point clouds and their corresponding BEV images. As can be seen, the point clouds in these datasets differ greatly, presenting sufficient challenges for evaluating reliable single-shot global localization. We detailed our setups of each dataset as follows:
1) KITTI [61]. This dataset contains a large number of point cloud data collected by a Velodyne 64-beam LiDAR. We select the sequences “00”, “02”, “05”, “06”, “07”, and “08” of the Odometry subset for evaluation since these sequences contain large revisited areas. We split the point clouds of each sequence into database frames and query frames for place retrieval. The partition of each sequence is summarized in Tab. II. We use the refined ground truth poses from semantic KITTI [67] with a distance threshold to determine if a loop closure exists.
TABLE II: Dataset Partition of the KITTI dataset.
Seq. 00 02 05 06 08
Db. 0-3000 0-3400 0-1000 0-600 0-3000
Query 3200-4650 3600-4661 1200-2751 800-1100 3200-4645
@@
4) RobotCar [65] and In-house [14]. These two datasets are broadly used by the recent place recognition method based on unordered points. The RobotCar dataset was created with a SICK LD-MRS LiDAR by repeatedly visiting a route of 10 km in Oxford. It contains 44 sequences collected on different days across a year. The In-house dataset consists of three scenarios including a university sector (U.S.), a residential area (R.A.), and a business district (B.D.). It is constructed from Velodyne-64 LiDAR scans and each of them contains have 5 sequences. Different from the aforementioned datasets that provide single LiDAR scans, these two datasets provide submaps built from consecutive scans.

### v3

@@ III BEV-based Global Localization Pipeline
Our method uses bird’s eye view (BEV) images as an intermediate representation of LiDAR data to perform global localization. As shown in Fig. 2, we project point clouds into BEV views and extract distinct local and global descriptors with the rotation equivariant and invariant network (REIN). Then, we employ a two-stage localization paradigm, i.e., place recognition followed by pose estimation. In place recognition, we perform place recognition with the global descriptors and find the top-matched BEV image from a pre-built database. In pose estimation, we exploit the local descriptors to estimate the relative pose between the query and matched BEV image pair using RANSAC. The global pose of the query is finally computed as the combination of the stored pose of the matched frame and the relative pose. In the following, we first explain the BEV image generation process, and then briefly describe the proposed REIN network. Finally, we elaborate on our proposed two-stage localization inference pipeline.
@@ III-A Bird’s Eye View Representation
Following existing 3-DoF localization works [31, 19, 57], we assume that when a ground vehicle moves within a small local area, it travels on a surface that is mostly flat. Based on this assumption, we generate BEV images through the orthogonal projection and concentrate on estimating the pose in 3-DoF, including (x, y, yaw). The (z, pitch, roll) can also be derived from the stored pose of the matched frame, but they are not the primary focus of this article. The BEV image representation shows benefits in many aspects of our hierarchical localization system. In place recognition, BEV images offer a comprehensive view of the distribution of road elements. Therefore, it is more intuitive and stable to extract global descriptors from such images to depict the structural information of the scene. In the context of pose estimation, the transformations of BEV image pairs are estimated by solving the BEV image matching problem, which is fast because of the lightweight BEV representation. The lightweight BEV representation also benefits the storage resource which is important for real-world applications and multi-robot communication.
@@
where $N_{g}$ denotes the number of points in the grid at position $(u,v)$ and $N_{m}$ the normalization factor. $N_{m}$ is set as the maximum value of the point cloud density.
Unlike traditional BEV image, also called elevation map [30, 31], which stores the maximum height of the points in each bin, we use the point densities. This is because the elevation map is influenced by the sensor’s orientation, as the maximum measured height changes with the distance between the scanner and the object. In contrast, the point density on a surface is less sensitive to viewpoint changes [51, 63].
@@ III-C Place Recognition
where $\mathbf{T}_{i}\in\mathbb{R}^{3\times 3}$ and $\mathbf{V}_{i}\in\mathbb{R}^{K\times C}$ are the corresponding global pose and global descriptor of the BEV image $\mathbf{I}_{i}$. The database construction is performed online or offline according to specific tasks. For localization within a global map, the database is typically constructed offline during the map-building process, far from the time of current application. For loop closure in SLAM, the database is constructed online, as loop closure operates as a SLAM module designed to detect loops and correct accumulated drift in real time.
2) Place recognition. For the query BEV image $\textbf{I}_{q}$, we generate its global descriptor $\textbf{V}_{q}$ with BEVPlace++ feature extractor. The matched frame is computed through the nearest global descriptor searching:
@@ III-D Pose Estimation
1) BEV image matching. We first extract FAST [64] keypoints from the BEV images to enable fast and accurate keypoint detection. Furthermore, on BEV images, the detected FAST keypoints usually have good repeatability since they are usually located at objects with vertical structures in the environment (e.g., poles, facades, guideposts). Moreover, the FAST detector requires no training, making it suitable for the weakly supervised design of our framework. We then assign each keypoint with a local descriptor interpolated from the REM feature map. We perform local feature matching between the BEV image pair and use RANSAC to estimate the relative transform with the keypoint correspondences.
@@ IV-A BEV Features Generation Through CNNs
We further plot the average feature distance with respect to translation displacements using different CNN backbones, including ResNet34 [65], VGG16 [66] and EfficientNet [67], on different datasets, such as KITTI [68], NCLT [69], and MCD [70], shown in Fig. 3 (e), (f), and (g). These datasets include point clouds collected by various types of LiDAR scanners with different fields of view and sparsity levels across diverse environments. As can be seen, spatially close features exhibit small feature distances while distant ones show larger feature distances, regardless of the dataset setup or used CNN backbones, which could be formulated as
@@
This inherent distinctiveness reveals that geometrically adjacent BEV points maintain adjacency in the feature space, highlighting the discriminative capability of the feature encoder and its ability to preserve spatial relationships. Such properties are crucial for accurately estimating pose, as they help in distinguishing between different keypoints and ensuring that their spatial relationships are preserved.
@@
Leveraging this characteristic, we establish BEV feature correspondences through nearest neighbor search and utilize RANSAC to solve for pose estimation.
Fig. 4 (a) illustrates feature matching results using ResNet34 under large translations. Each row corresponds to BEV images from the KITTI, NCLT, and MCD datasets, respectively. Following pose estimation, the overlaid BEV image pairs indicate our method’s capability to accurately estimate significant translation movements. It is noted that feature matching could fail under rotations, as shown in Fig. 4 (b), since traditional CNNs are not rotation-equivariant and the feature distinctiveness is only preserved under translation transformations.
@@ IV-B Rotation Equivariant Module
As discussed above, traditional CNNs are translation equivariant but cannot handle rotations well. To achieve more robust pose estimations, we further design a network to extract rotation-equivariant feature maps from BEV images. We introduce a simple and effective rotation equivariant module (REM) to extract rotation equivariant local features. REM uses modern CNNs as basic modules to keep their representation ability to BEV images and achieves rotation equivariance by using the maximum CNN feature response of BEV images under different rotation transforms. It is noted that there are some rotation equivariant networks that have been used in vision tasks such as object detection [61, 60]. Our REM differs from them since REM uses CNN backbone as basic modules and thus is explicitly designed to retain the descriptive capability of CNN backbones to BEV images.
@@
Figure 6: Comparison between ResNet and REM features with rotations. Features are extracted from a pair of rotated BEV images with ResNet and REM. The difference between the ResNet features of the rotated BEV and the rotated ResNet features of the original BEV is large. On the other hand, REM shows rotation equivariance, and the difference between the REM features of the rotated BEV and the rotated REM features of the original BEV tends to be zero.
@@
$=\max_{r\in\mathcal{R}}\textbf{R}_{\alpha}\textbf{R}_{\alpha}^{-1}\textbf{R}^{-1}_{r}\circ{\phi((\textbf{R}_{r}\textbf{R}_{\alpha})\circ\textbf{I})}$
$=\max_{r\in\mathcal{R}}\textbf{R}_{\alpha}\textbf{R}_{r+\alpha}^{-1}\circ{\phi((\textbf{R}_{r+\alpha})\circ\textbf{I})}$
Since the rotation group $\mathcal{R}$ is closed under addition, the mapping $r\mapsto r+\alpha$ is a bijection on $\mathcal{R}$. Thus, the set $\{\textbf{R}_{r+\alpha}^{-1}\circ\phi(\textbf{R}_{r+\alpha}\circ\textbf{I})|r\in\mathcal{R}\}$ is identical to the set $\{\textbf{R}_{r}^{-1}\circ\phi(\textbf{R}_{r}\circ\textbf{I})|r\in\mathcal{R}\}$. Applying a left multiplication by an additional rotation $R_{\alpha}\in\mathcal{R}$ to each element of the set does not alter the identity. Thus, we can rewrite $\textbf{F}^{\prime}$ as:
$\textbf{F}^{\prime}$ $=\max_{r\in\mathcal{R}}\textbf{R}_{\alpha}\textbf{R}_{r+\alpha}^{-1}\circ{\phi((\textbf{R}_{r+\alpha})\circ\textbf{I})}$ (13)
$=\max_{r\in\mathcal{R}}\textbf{R}_{\alpha}\textbf{R}_{r}^{-1}\circ{\phi(\textbf{R}_{r}\circ\textbf{I})}.$
The maximum of a set is determined solely by the largest element within it. Therefore, the result remains the same regardless of the order of the elements in the set. Additionally, rotation operations only induce spatial permutations without altering the values (ignoring edge effects), we have:
$\textbf{F}^{\prime}$ $=\max_{r\in\mathcal{R}}\textbf{R}_{\alpha}\textbf{R}_{r}^{-1}\circ{\phi(\textbf{R}_{r}\circ\textbf{I})}$ (14)
$=\textbf{R}_{\alpha}\circ\max_{r\in\mathcal{R}}\textbf{R}_{r}^{-1}\circ{\phi(\textbf{R}_{r}\circ\textbf{I})}$
@@
This derivation demonstrates that the output features F are rotation-equivariant, as the effect of a rotation $\textbf{R}_{\alpha}$ on the input image I results in the same rotation applied to the resulting feature map F. Theoretically, we need to sample infinite angles to achieve continuous rotation equivariance. In our experiments, we find that a small $N$ can achieve satisfactory performance. This is possible because the pooling and downsample operations in CNN could provide some robustness to small movements.
@@
$\mathbf{V}_{k}=\sum_{i}a_{k}(\mathbf{f}_{i})(\mathbf{f}_{i}-\mathbf{c}_{k}),$ (16)
where $a_{k}(\mathbf{f}_{i})$ the soft-assignment of the feature $\mathbf{f}_{i}$ for the $k$-th cluster center, computed as
$a_{k}(\mathbf{f}_{i})=\frac{e^{\mathbf{w}_{k}^{T}\mathbf{f}_{i}+b_{k}}}{\sum_{k^{\prime}}e^{{\mathbf{w}_{k^{\prime}}^{T}\mathbf{f}_{i}+b_{k^{\prime}}}}}$ (17)
2) Rotation Invariance Analysis. Let $\{\mathbf{f}_{i}\}_{i=1,...,{H^{\prime}W^{\prime}}}$ and $\{\mathbf{f^{\prime}}_{i}\}_{i=1,...,{H^{\prime}W^{\prime}}}$ denote the local descriptor sets of the raw and the rotated BEV image. As the REM feature map is rotation equivariant, set $\{\mathbf{f}_{i}\}_{i=1,...,H^{\prime}W^{\prime}}$ and $\{\mathbf{f^{\prime}}_{i}\}_{i=1,...,H^{\prime}W^{\prime}}$ contain the same elements but differ in their permutation due to the rotation. This indicates a one-to-one correspondence between the two sets, ensuring consistency in their cluster assignments. Since summation $\sum$ is a symmetric operation with respect to permutations, we have
$\mathbf{V}_{k}^{\prime}$ $=a_{k}(\mathbf{f}_{1}^{\prime})(\mathbf{f}_{1}^{\prime}-\mathbf{c}_{k})+a_{k}(\mathbf{f}_{2}^{\prime})(\mathbf{f}_{2}^{\prime}-\mathbf{c}_{k})+...+a_{k}(\mathbf{f}_{n}^{\prime})(\mathbf{f}_{n}^{\prime}-\mathbf{c}_{k})$ (18)
$=\sum_{i}a_{k}(\mathbf{f}_{i}^{\prime})(\mathbf{f}_{i}^{\prime}-\mathbf{c}_{k})$
$=\sum_{i}a_{k}(\mathbf{f}_{i})(\mathbf{f}_{i}-\mathbf{c}_{k})=\mathbf{V}_{k}.$
This equality demonstrates that the global descriptors $\mathbf{V}_{k}$ are invariant to rotation, as they include symmetric operations that are unaffected by the permutation of local descriptors.
@@
Dataset KITTI [68] NCLT  [69] MCD [70] RobotCar [72] Inhouse [14] Apollo [73] Mulran [74] Our UGV
LiDAR Velodyne HDL-64E Velodyne HDL-32E Livox Mid-70 SICK LD-MRS 3D Velodyne HDL-64 Velodyne HDL-64 Ouster OS1-64 RoboSense Helios-32
FoV (HxV) ${}^{\circ}\times$ 26.8 ${}^{\circ}\times$ 41.3 ${}^{\circ}\times$ 70 ${}^{\circ}\times$ 3.2 ${}^{\circ}\times$ 26.8 ${}^{\circ}\times$ 26.8 ${}^{\circ}\times$ 45.0 ${}^{\circ}\times$ 70
# of Seq. 5 7 6 44 15 6 2 1
Scenes Country, Highway Campus City City Campus City, Highway City Campus
Cities Karlsruhe Michigan Singapore Oxford Singapore California Sejong Changsha
Traj. length 5 km 10 km 2km 10 km 10 km 73 km 23 km 1.5 km
Time span Single days Across 1 year Several days Across 1 year Several days Several days Several days Several days
Tasks PR & LC & GL PR & LC PR & GL PR PR GL GL LC
@@ V-A Datasets
We use 7 public datasets to thoroughly evaluate the performance of our method across various LiDAR sensor setups and data collected from different cities, including KITTI [68], NCLT [69], MCD [70], RobotCar [72], In-house [14], Apollo [73], and Mulran [74] datasets. We evaluate loop closing performance on KITTI and NCLT datasets since their data sequences have large loops. We test global localization on KITTI, NCLT, MCD, Apollo, and Mulran datasets as they have accurate ground truth poses. We evaluate place recognition on all the datasets. Our evaluation datasets cover diverse scenes, including city, countryside, and campus, and are collected in large-scale places under large time spans. The point cloud data are of different sparsity and different fields of view due to the usage of the various types of LiDAR sensors. Tab. I summarizes the meta information of the datasets. The point clouds in these datasets differ greatly, presenting sufficient challenges for evaluating reliable single-shot global localization. We further deploy our method on a real UGV platform equipped with a RoboSense-32 LiDAR (See Sec. VI-E). We detailed setups of each dataset as follows:
1) KITTI [68]. This dataset contains a large number of point cloud data collected by a Velodyne 64-beam LiDAR. We select the sequences “00”, “02”, “05”, “06”, “07”, and “08” of the Odometry subset for evaluation since these sequences contain large revisited areas. We split the point clouds of each sequence into database frames and query frames for place recognition, following the widely adopted partition [52, 17, 18]. We use the refined poses from semantic KITTI [76] with a distance threshold to determine if a loop closure exists.
@@
4) RobotCar [72] and In-house [14]. Both datasets are broadly used by the recent place recognition method based on unordered points. The RobotCar dataset was created with a SICK LD-MRS LiDAR by repeatedly visiting a route of 10 km in Oxford. It contains 44 sequences collected on different days across a year. The In-house dataset consists of three scenarios, including a university sector (U.S.), a residential area (R.A.), and a business district (B.D.). It is constructed from Velodyne-64 LiDAR scans, and each contains 5 sequences. Different from the aforementioned datasets that provide single LiDAR scans, these two datasets provide submaps built from consecutive scans.
5) Apollo [73]. This dataset is collected through multiple traversals along six different routes in the southern San Francisco Bay Area. It encompasses a diverse range of environments, including residential neighborhoods, urban downtown areas, and highways. The data is captured using a Velodyne HDL 64E rotating LiDAR, which provides a range and number of scanned 3D points comparable to the MulRan dataset. All six sequences are used for evaluation.
6) Mulran [74]. This dataset contains point clouds collected in South Korea using Ouster OS1-64 rotating LiDAR. It provides accurate ground truth generated by VRS-GPS/INS integrated navigation system and a graph SLAM. Following EgoNN [16], we use the longest and most diverse sequences collected in Sejong as the evaluation sequences.